# Estimating Cognitive Effort from Functional Near-Infrared Spectroscopy (fNIRS) Signals using Machine Learning

### v1

@@ Abstract.
Machine learning, cognitive effort, relative neural efficiency, relative neural involvement, performance score, functional Near-Infrared Spectroscopy (fNIRS), brain signal, hemodynamic response, educational games
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: XXXXX; XX; isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
@@ 1. Introduction
This is where the idea of cognitive effort becomes useful. Unlike cognitive load, which only shows how much the brain is working, cognitive effort considers both brain activation and performance together (Getchell and Shewokis, 2023; Koiler et al., 2022). When both brain activity is high and performance is also good, we interpret that as involvement. But if the brain is working hard and performance is still poor, it likely reflects burden. In other words, CE helps us differentiate between productive effort and mental overload. Even when performance looks efficient, CE can reveal whether that efficiency required excessive mental strain. By combining brain load and performance, we get a clearer picture of whether the effort was manageable or overwhelming(Getchell and Shewokis, 2023; Koiler et al., 2022).
@@ 3.1.1. Apparatus
Figure 2 presents study setup and fNIRS headband orientation. We used two computers and an fNIRS device (Figure 2 (a)). We used an 18-channel headband fNIRS imager 2000S (fNIRS Device LLC, Potomac, MD, USA), 4 light emitters, wavelength ($\lambda$) 730-850 nm with average power $<1$ mW) sampled at 10 Hz. The detectors were separated by 2.5 cm with a penetrating depth of approximately 1.2 cm. This headband was placed on participants’ foreheads, and fNIRS collected data for 16 channels (Figure 2(b)). Channel 1-4 and channel 13-16 collectively called Lateral PFC (LPFC) and Channel 5-12 called Ventromedial PFC (VMPFC). Participants interacted with the educational quiz on a Dell XPS touchscreen laptop (L1) [Intel (R) Core(TM) i7-7700HQ processor at 2.80 GHz]. Another desktop computer (L2) [Intel(R) Core(TM) i7-10700T processor at 2.00 GHz] was connected to the fNIRS device. The laptop was used to collect fNIRS data using Cognitive Optical Brain Imaging (COBI) Studio Software. We also used fNIRSoft Software (Version 4.9) to apply filters on $\Delta HbO$ data and generate CSV file for further processing.
@@ 3.1.2. Study Procedure
After completing a ten-question pre-test, they watched the recorded video tutorial where they learned the basic terminologies of the graph. In our study, we divided our 16 questions into two sessions, with each session having two segments and each segment having four questions. Participants had up to 30 seconds to answer each question, and after each question, they had five seconds of feedback. So, each segment was 140 seconds $((4\times 30=120+4\times 5=20)=140)$ long. In between the two segments, there was a 20-second rest period. After the first session, they had a 6-10 minute break. After each session, the participants completed a post-test. We randomized the question order to minimize learning effects and control for difficulty bias across participants.
@@ 3.1.3. Cognitive Effort Measure
[…]
@@ 3.2. Data Acquisition
We collected fNIRS data using COBI, and the fNIRS recording rate was 10 samples per second. The data was saved in fNIRS format. Each question was presented for 30 seconds, so each file contained 300 rows for each question corresponding to 16 optode ($\Delta HbO$) values. Sixteen participants answered 16 questions each, leading to 256 responses in the dataset ($16\times 16$). The dataset has 168 correct responses (class 1) and 88 incorrect responses (class 0). We used only $\Delta HbO$ because previous studies showed that $\Delta HbO$ signals are more sensitive to changes in blood flow in the participant’ cortex relative to $\Delta HbR$ signals (Gui et al., 2025; Liu et al., 2019; Pan et al., 2018). Since participants often answered within 20 seconds, only the first 200 rows per question were retained. The dataset contains 819,200 data points:
16 participants×16 questions×200 $16\text{ participants}\times 16\text{ questions}\times 200\text{ fNIRS}_{\text{points}}\times 16\text{ optodes}$.
@@ 3.4. Feature Extraction Framework
Statistical Features (ST): A total of $128$ statistical features were extracted from $\Delta HbO$ signals, computed as $16$ optodes $\times$ $8$ statistical measures (mean, standard deviation, maximum, minimum, gradient mean, squared gradient mean, skewness, and kurtosis).
Functional Connectivity (FC) Features: For each question, we computed a $16\times 16$ Pearson correlation matrix using the 200-timepoint $\Delta HbO$ signals from all 16 optodes. These pairwise correlations capture functional connectivity between brain regions. To avoid redundancy, we extracted the 120 unique values from the upper triangle of the matrix (excluding the diagonal) as FC features.
@@
3.5. Model Implementation and Training Details
All models were implemented using scikit-learn (v1.4) and XGBoost (v2.0) in Python. Before training the models, we applied standardization (Z-score normalization) to ensure uniform feature scaling Finally, we used 5-fold cross-validation by applying GroupKFold $(n\_splits=5)$ to ensured participant-independent splits. Each fold consisted of a participant-wise split to ensure no overlap between train and test participants. Specifically, in each fold, approximately 80% of participants (13 or 12 out of 16) were used for training and the remaining 20% (3 or 4) for testing. This setup ensured participant-independent evaluation and avoided data leakage across folds. For preprocessing, all features were standardized using z-scores via StandardScaler.
The models and their key hyperparameters were as follows:
  • Logistic Regression: max_iter=1000
  • SVM (RBF): kernel=’rbf’, probability=True
  • KNN: n_neighbors=5
  • LDA: default parameters
  • Decision Tree: random_state=42
  • Random Forest: n_estimators=100
  • XGBoost: use_label_encoder=False, eval_metric=’logloss’
All models were evaluated using accuracy, precision, recall, and F1 score averaged across the 5 folds. Precision measures the proportion of correctly predicted positive observations to the total predicted positive observations. Recall evaluates the proportion of correctly predicted positive observations to all actual positive observations. A high recall indicates the model successfully captures most of the positive instances. F1-score is the harmonic mean of precision and recall. It is particularly useful when the class distribution is imbalanced.
To evaluate whether ML can reliably predict performance from fNIRS-derived features, we assessed multiple classifiers (e.g., Logistic Regression, SVM, Random Forest) across five types of features (Basic, statistical features (ST), Functional Connectivity (FC), ST+FC, and Temporal). These features are then used as input for various classification algorithms. Five different classifiers are used independently within a 5-fold stratified cross-validation framework: Logistic Regression (LR), SVM (RBF Kernel), K-Nearest Neighbors (KNN, k=5), Linear discriminant analysis (LDA), Decision Tree (DT), Random Forest (RF), XGBoost. The output of the classification is a binary prediction of correct (1) or incorrect (0) response.
@@ 6. Conclusion
Acknowledgments
We wish to express our gratitude to the study participants and lab members.

### v2

# Beyond Load: Understanding Cognitive Effort through Neural Efficiency and Involvement using fNIRS and Machine Learning

@@ Abstract.
Cognitive effort, cognitive load, relative neural efficiency, relative neural involvement, performance score, functional Near-Infrared Spectroscopy (fNIRS), brain signal, hemodynamic response, educational games, machine learning,
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: CHI; 13–17 April, 2026; Barcelonaisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
@@ 1. Introduction
This is where the idea of cognitive effort becomes useful. Unlike cognitive load, which only shows how much the brain is working, cognitive effort considers both brain activation and performance together (Getchell and Shewokis, 2023; Koiler et al., 2022; Sharmin et al., 2024c, 2025). When both brain activity is high and performance is also good, we interpret that as involvement. But if the brain is working hard and performance is still poor, it likely reflects burden. In other words, CE helps us differentiate between productive effort and mental overload. Even when performance looks efficient, CE can reveal whether that efficiency required excessive mental strain. By combining brain load and performance, we get a clearer picture of whether the effort was manageable or overwhelming(Sharmin et al., 2025, 2024c; Getchell and Shewokis, 2023; Koiler et al., 2022).
@@ 3.1.1. Apparatus
Figure 2 presents study setup and fNIRS headband orientation. We used two computers and an fNIRS device (Figure 2 (a)). We used an 18-channel headband fNIRS imager 2000S (fNIRS Device LLC, Potomac, MD, USA), 4 light emitters, wavelength ($\lambda$) 730-850 nm with average power <1<1< 1 mW) sampled at 10 Hz. The detectors were separated by 2.5 cm with a penetrating depth of approximately 1.2 cm. This headband was placed on participants’ foreheads, and fNIRS collected data for 16 channels (Figure 2(b)). Channel 1-4 and channel 13-16 collectively called Lateral PFC (LPFC) and Channel 5-12 called Ventromedial PFC (VMPFC). Participants interacted with the educational quiz on a Dell XPS touchscreen laptop (L1) [Intel (R) Core(TM) i7-7700HQ processor at 2.80 GHz]. Another desktop computer (L2) [Intel(R) Core(TM) i7-10700T processor at 2.00 GHz] was connected to the fNIRS device. The laptop was used to collect fNIRS data using Cognitive Optical Brain Imaging (COBI) Studio Software. We also used fNIRSoft Software (Version 4.9) to apply filters on $\Delta HbO$ data and generate CSV file for further processing.
@@ 3.1.2. Study Procedure
After completing a ten-question pre-test, they watched the recorded video tutorial where they learned the basic terminologies of the graph. In our study, we divided our 16 questions into two sessions, with each session having two segments and each segment having four questions. Participants had up to 30 seconds to answer each question, and after each question, they had five seconds of feedback. So, each segment was 140 seconds ((4×30=120+4×5=20)=140)((4$\times$ 30=120+4$\times$ 5=20)=140)( ( 4 × 30 = 120 + 4 × 5 = 20 ) = 140 ) long. In between the two segments, there was a 20-second rest period. After the first session, they had a 6-10 minute break. After each session, the participants completed a post-test. We randomized the question order to minimize learning effects and control for difficulty bias across participants.
@@ 3.1.3. Cognitive Effort Measure
(2) $\hskip 56.9055ptCE_{z}=\dfrac{\dfrac{1}{\Delta HbO_{i}}-\dfrac{1}{\Delta HbO_{\text{(GM)}}}}{\dfrac{1}{\Delta HbO_{\text{(SD)}}}}\hskip 28.45274pt$
@@ 3.2. Data Acquisition
We collected fNIRS data using COBI, and the fNIRS recording rate was 10 samples per second. The data was saved in fNIRS format. Each question was presented for 30 seconds, so each file contained 300 rows for each question corresponding to 16 optode ($\Delta HbO$) values. Sixteen participants answered 16 questions each, leading to 256 responses in the dataset ($\times 16$ 16). The dataset has 168 correct responses (class 1) and 88 incorrect responses (class 0). We used only $\Delta HbO$ because previous studies showed that $\Delta HbO$ signals are more sensitive to changes in blood flow in the participant’ cortex relative to $\Delta HbR$ signals (Gui et al., 2025; Liu et al., 2019; Pan et al., 2018). Since participants often answered within 20 seconds, only the first 200 rows per question were retained. The dataset contains 819,200 data points:
$16\text{ participants}\times 16\text{ questions}\times 200\text{ fNIRS}_{\text{points}}\times 16\text{ optodes}$.
@@ 3.4. Feature Extraction Framework
Statistical Features (ST): 16×8=12816$\times$ 8=12816 × 8 = 128 statistical features (ST) were derived per optode which includes mean, standard deviation, maximum, minimum, gradient mean, squared gradient mean, skewness, and kurtosis.
Functional Connectivity (FC) Features: For each question, we computed a $\times 16$ 16 Pearson correlation matrix using the 200-timepoint $\Delta HbO$ signals from all 16 optodes. These pairwise correlations capture functional connectivity between brain regions. To avoid redundancy, we extracted the 120 unique values from the upper triangle of the matrix (excluding the diagonal) as FC features.
@@
3.5. Model Architecture
To evaluate our model, we compute precision, recall, and F1-score. Precision measures the proportion of correctly predicted positive observations to the total predicted positive observations. Recall evaluates the proportion of correctly predicted positive observations to all actual positive observations. A high recall indicates the model successfully captures most of the positive instances. F1-score is the harmonic mean of precision and recall. It is particularly useful when the class distribution is imbalanced. To evaluate whether ML can reliably predict performance from fNIRS-derived features, we assessed multiple classifiers (e.g., Logistic Regression, SVM, Random Forest) across five types of features (Basic, statistical features (ST), Functional Connectivity (FC), ST+FC, and Temporal). Before training the models, we applied standardization (Z-score normalization) to ensure uniform feature scaling Finally, we used 5-fold cross-validation by applying GroupKFold $(n\_splits=5)$ to ensured participant-independent splits. Each fold consisted of a participant-wise split to ensure no overlap between train and test participants. Specifically, in each fold, approximately 80% of participants (13 or 12 out of 16) were used for training and the remaining 20% (3 or 4) for testing. This setup ensured participant-independent evaluation and avoided data leakage across folds.
These features are then used as input for various classification algorithms. Five different classifiers are used independently within a 5-fold stratified cross-validation framework: Logistic Regression (LR), SVM (RBF Kernel), K-Nearest Neighbors (KNN, k=5), Linear discriminant analysis (LDA), Decision Tree (DT), Random Forest (RF), XGBoost. The output of the classification is a binary prediction of correct (1) or incorrect (0) response.